Borax, continued:


An aqueous solution of Borax (Na2B4O7) is sometimes used to fireproof wood. 


How many moles of Na2B4O7 are in 600.0 ml of a 1.57 M Na2B4O7 solution? 


1
Expert's answer
2022-03-31T01:51:04-0400

Volume of solution = 600 mL


Molarity = 1.57


"Molarity = \frac{mole}{volume}\times 1000"


"1.57 = \frac{mole}{600}\times 1000"


mole= 0.942